The left and right page numbers of an open book are two consecutive integers whose sum is 541.Find these page numbers.

1
Expert's answer
2020-08-27T04:24:52-0400

Let "x" be the left page number, then "x+1" is the right page number.

We have "x+(x+1)=541", that is "2x+1=541". So "x=\frac{541-1}{2}=270" and "x+1=271"

Answer: "270, 271"